# New criteria for the rectifiability of Radon measures in terms of Riesz transforms

### v1

@@ Abstract.
If $\delta_{1}$ is small enough, depending on $n$ and $\alpha$, and $\varepsilon$} is small enough, then there exists some uniformly $n$-rectifiable set $\Gamma$ and some $\tau>0$ such that $\mu(\Gamma\cap B_{0})\geq\tau\,\mu(B_{0}).$
@@ 1. Introduction
We recall that a Borel measure $\mu$ in ${\mathbb{R}}^{d}$ is called $n$-rectifiable if there is a countable collection of Lipschitz maps $g_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{d}$ such that $\mu({\mathbb{R}}^{d}\setminus\bigcup_{i}g_{i}({\mathbb{R}}^{n}))=0$ and moreover $\mu$ is absolutely continuous with respect to the Hausdorff measure ${\mathcal{H}}^{n}$. A set $E\subset{\mathbb{R}}^{d}$ is $n$-rectifiable if the Hausdorff measure ${\mathcal{H}}^{n}|_{E}$ is $n$-rectifiable. On the other hand, $E$ is called purely $n$-unrectifiable if ${\mathcal{H}}^{(}E\cap F)=0$ for any $n$-rectifiable set $F\subset{\mathbb{R}}^{d}$.
@@
The next corollary deals with reflectionless measures. For our purposes, a reflectionless measure for the $n$-dimensional Riesz transform is a Radon measure $\mu$ in ${\mathbb{R}}^{d}$ having $n$-polynomial growth (i.e, satisfying (1.1)), such that ${\mathcal{R}}_{\mu}$ is bounded in $L^{2}(\mu)$, and such that ${\mathcal{R}}\mu$ is zero in a $\mathop{\mathrm{BMO}}(\mu)$ sense, i.e., ${\mathcal{R}}\mu$ is constant $\mu$-a.e. In [JN2], Jaye and Nazarov have shown that if all the $n$-Ahlfors reflectionless measures for the Riesz transform are $n$-flat, i.e., of the form $c\,{\mathcal{H}}^{n}|_{L}$ for some $n$-plane $L$, then the the $n$-dimensional David-Semmes problem has a positive answer. That is, for any $n$-Ahlfors regular set $E\subset{\mathbb{R}}^{d}$, the $L^{2}({\mathcal{H}}^{n}|_{E})$ boundedness of ${\mathcal{R}}_{{\mathcal{H}}^{n}|_{E}}$ implies the uniform $n$-rectifiability of $E$. This question is still open in ${\mathbb{R}}^{d}$ when $1<n<d$. However, up to now the structure of reflectionless measure for the Riesz transform is not well understood. It is not known if they are all $n$-flat even in the $n$-Ahlfors regular case with n=d1n=d-1. For more information on reflectionless measures, see also [JN1].
We remark that by now it is not known if there exist reflectionless measures $\mu$ such that the set $\{x\in\operatorname{supp}(\mu):\theta_{\mu}^{n,*}(x)=0\big\}$ has positive $\mu$-measure. In this paper we will prove the following.
@@
We also remark that in Section 8.4 we show a somewhat weaker variant of Theorem 1.5 which suffices for some applications to harmonic measure. This variant has the advantage that its proof does not rely on the lengthy arguments from [DT]. Concerning the applications to harmonic measure, we mention that in a forthcoming work of Luis Lloret and the author [LT], the following result is obtained by using Theorem 1.5 as an essential tool.
@@
Remark that, by the definition, $\Theta_{\omega}(x,r)\to 0$ as $r\to 0$ for every $x\in Z$ with $Z\subset{\partial\Omega}$ as in the theorem. Then (1.4) asserts that $\Theta_{\omega}(x,r)$ has arbitrarily large relative oscillations as it tends to 0 (equivalently, the density $\Theta_{\omega}(x,\cdot)$ fails to be almost-increasing near $r=0$). We recall that a deep result due to Wolff [Wo1], in the complex plane, asserts that for any open set $\Omega\subset{\mathbb{R}}^{2}$, harmonic measure is concentrated in a set of $\sigma$-finite length, which implies that $\omega(Z)=0$ in this case. In higher dimensions this is not the case, and in fact, one may have $\dim_{H}(\omega)>n$ for $\Omega\subset{\mathbb{R}}^{n+1}$, by another celebrated result of Wolff [Wo2].
@@ 3. The approximation lemma
$|{{\widetilde{{\mathcal{R}}}}_{\varepsilon}}\mu(x)-{{\widetilde{{\mathcal{R}}}}_{\varepsilon}}\mu(y)|\lesssim\int|K_{\varepsilon}(x-z)-K_{\varepsilon}(x-y)|\,d\mu(y)\lesssim\int\frac{\ell_{k}}{({\varepsilon}+|x-z|)^{n+1}}\,d\mu(z)\lesssim{\mathcal{P}}_{\mu}(Q).$
@@
$\int_{2B_{0}}|{\mathcal{R}}\sigma_{k}-m_{\sigma_{k},2B_{0}}({\mathcal{R}}\sigma_{k})|^{2}\,d\sigma_{k}$ $\leq\int_{2B_{0}}|{\mathcal{R}}\sigma_{k}(x)-\rho|^{2}\,d\sigma_{k}$
$\leq 2\int_{2B_{0}}|{\widetilde{{\mathcal{R}}}}_{\ell_{k}}\mu(x)-\rho|^{2}\,d\mu+C\int_{2B_{0}}{\mathcal{P}}_{\mu}(B(x,\ell_{k}))^{2}\,d\mu(x).$
@@
$\lim_{k\to\infty}\int_{2B_{0}}|{\widetilde{{\mathcal{R}}}}_{\ell_{k}}\mu(x)-\rho|^{2}\,d\mu=\int_{2B_{0}}|{\mathcal{R}}\mu(x)-\rho|^{2}\,d\mu=\int_{2B_{0}}|{\mathcal{R}}\mu(x)-m_{\mu,2B_{0}}({\mathcal{R}}\mu)|^{2}\,d\mu.$
@@ 5. The main lemma
Hence there is a compact set $E\subset B$ such that $\mu(E)\geq\mu(B)/4$ and $M_{n}(\chi_{B}\mu)\leq 2C_{2}\,\Theta_{\mu}(B)$ for all $x\in E$. So the measure
@@
    (5.3) $|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))(x)\leq 16\lambda\,\Theta_{\mu}(B_{1})^{p}.$
@@
The fact that $F(a)\leq F(1)\leq 4\lambda\sigma_{p}(B_{0})$ also implies that
@@
To prove (c), for any ball $B$ centered in $R_{0}\cap\operatorname{supp}\mu$ and any $t\in(0,1)$, we consider the function
@@
Notice that $g$ is differentiable in $[0,1)$ and $g(t)\leq F(a_{t})\leq F(a)=g(0)$, by the minimizing property of $a$. Hence $g^{\prime}(0+)\geq 0$. Computing, we get
@@
Let $\Omega\subset{\mathbb{R}}^{n+1}$ be a bounded open set, let $\xi\in\partial\Omega$, and let $r_{0}>0$ and $r_{1}=8^{-N}r_{0}$, for some $N\geq 1$. Let $u$ be a nonnegative function which is continuous in $\overline{B(\xi,r)\cap\Omega}$ and subharmonic in $B(\xi,r)\cap\Omega$, and vanishes on $B(\xi,r)\cap\partial\Omega$. Suppose that there are $c_{3}>0$ and a subset $J\subset\{1,\ldots,N\}$ with $J\geq\kappa N$, for some $\kappa>0$, such that
@@
Let $\mu$ and $B_{0}$ satisfy the assumptions of Lemma 5.1 and let $R_{0}$, $c_{R_{0}}$ be as in Lemma 5.3. Suppose that, for some $p\in(1,1+\frac{1}{n})$ and some $\lambda\in(0,1)$,
@@
$|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))(x)\leq\big[16\,\lambda+C(\alpha,\eta,\tau)\,\delta_{1}^{\kappa_{1}}\big]\,\Theta_{\mu}(B_{1})^{p}$
@@
$f(x)=\max\big(0,\,|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))(x)-16\lambda\,\Theta_{\mu}(B_{1})^{p}\big).$
@@
$\quad+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))(x).$
@@
(5.13) $|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))(x)\leq\gamma\,\Theta_{\mu}(B_{1})^{p}$
@@
$\leq C\gamma\Theta_{\mu}(B_{1})^{p}\int|\nabla{\varphi}|\,d{\mathcal{L}}^{n+1}-p\int{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))\,|\nabla{\varphi}|\,d{\mathcal{L}}^{n+1}$
$\leq C\gamma\Theta_{\mu}(B_{1})^{p}\,\mu(B_{1})-p\int_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))\,{\mathcal{R}}_{{\mathcal{L}}^{n+1}}(|\nabla{\varphi}|)\,d\nu.$
To bound the last integral on the right hand side we use the fact that, for any $x\in{\mathbb{R}}^{d}$,
@@
$\bigg|\int_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}$ $|^{p-2}({\mathcal{R}}\nu(x)-c_{R_{0}}))\,{\mathcal{R}}_{{\mathcal{L}}^{n+1}}(|\nabla{\varphi}|)\,d\nu\bigg|$
$\lesssim\Theta_{\mu}(B_{1})\int_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p-1}\,d\nu$
$\leq\Theta_{\mu}(B_{1})\left(\int_{R_{0}}|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}\,d\nu\right)^{1/p^{\prime}}\nu(R_{0})^{1/p}$

### v2

@@ Abstract.
If $\delta_{1}$ is small enough, depending on $n$ and $\alpha$, and $\varepsilon$} is small enough, then there exists some uniformly $n$-rectifiable set $\Gamma$ and some $\tau>0$ such that $\mu(\Gamma\cap 2B_{0})\geq\tau\,\mu(B_{0}).$
@@ 1. Introduction
We recall that a Borel measure $\mu$ in ${\mathbb{R}}^{d}$ is called $n$-rectifiable if there is a countable collection of Lipschitz maps $g_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{d}$ such that $\mu({\mathbb{R}}^{d}\setminus\bigcup_{i}g_{i}({\mathbb{R}}^{n}))=0$ and moreover $\mu$ is absolutely continuous with respect to the Hausdorff measure ${\mathcal{H}}^{n}$. A set $E\subset{\mathbb{R}}^{d}$ is $n$-rectifiable if the Hausdorff measure ${\mathcal{H}}^{n}|_{E}$ is $n$-rectifiable. On the other hand, $E$ is called purely $n$-unrectifiable if ${\mathcal{H}}^{n}(E\cap F)=0$ for any $n$-rectifiable set $F\subset{\mathbb{R}}^{d}$.
@@
The next corollary deals with reflectionless measures. For our purposes, a reflectionless measure for the $n$-dimensional Riesz transform is a Radon measure $\mu$ in ${\mathbb{R}}^{d}$ having $n$-polynomial growth (i.e, satisfying (1.1)), such that ${\mathcal{R}}_{\mu}$ is bounded in $L^{2}(\mu)$, and such that ${\mathcal{R}}\mu$ is zero in a $\mathop{\mathrm{BMO}}(\mu)$ sense, i.e., ${\mathcal{R}}\mu$ is constant $\mu$-a.e. In [JN2], Jaye and Nazarov have shown that if all the $n$-Ahlfors reflectionless measures for the Riesz transform are $n$-flat, i.e., of the form $c\,{\mathcal{H}}^{n}|_{L}$ for some $n$-plane $L$, then the the $n$-dimensional David-Semmes problem has a positive answer. That is, for any $n$-Ahlfors regular set $E\subset{\mathbb{R}}^{d}$, the $L^{2}({\mathcal{H}}^{n}|_{E})$ boundedness of ${\mathcal{R}}_{{\mathcal{H}}^{n}|_{E}}$ implies the uniform $n$-rectifiability of $E$, which is still an open question in ${\mathbb{R}}^{d}$ for $1<n<d$. However, up to now the structure of reflectionless measure for the Riesz transform is not well understood. It is not known if they are all $n$-flat even in the $n$-Ahlfors regular case with n=d1n=d-1. For more information on reflectionless measures, see also [JN1].
We also remark that by now it is not known if there exist reflectionless measures $\mu$ such that the set $\{x\in\operatorname{supp}(\mu):\theta_{\mu}^{n,*}(x)=0\big\}$ has positive $\mu$-measure. In this paper we will prove the following.
@@
We also remark that in Section 8.4 we show a somewhat weaker variant of Theorem 1.5 which suffices for some applications to harmonic measure. This variant has the advantage that its proof does not rely on the lengthy arguments from [DT].
Concerning the applications to harmonic measure, we mention that in a forthcoming work of Luis Lloret and the author [LT], the following result is obtained by using Theorem 1.5 as an essential tool.
@@
Remark that, by the definition, $\Theta_{\omega}(x,r)\to 0$ as $r\to 0$ for every $x\in Z$ with $Z\subset{\partial\Omega}$ as in the theorem. In a sense, (1.4) asserts that $\Theta_{\omega}(x,r)$ has arbitrarily large relative oscillations as it tends to 0 (equivalently, the density $\Theta_{\omega}(x,\cdot)$ fails to be almost-increasing near $r=0$). We recall that a deep result due to Wolff [Wo1], in the complex plane, asserts that for any open set $\Omega\subset{\mathbb{R}}^{2}$, harmonic measure is concentrated in a set of $\sigma$-finite length, which implies that $\omega(Z)=0$ in this case. In higher dimensions this is not the case, and in fact, one may have $\dim_{H}(\omega)>n$ for $\Omega\subset{\mathbb{R}}^{n+1}$, by another celebrated result of Wolff [Wo2].
@@ 3. The approximation lemma
$|{{\widetilde{{\mathcal{R}}}}_{\varepsilon}}\mu(x)-{{\widetilde{{\mathcal{R}}}}_{\varepsilon}}\mu(y)|\lesssim\int|K_{\varepsilon}(x-z)-K_{\varepsilon}(y-z)|\,d\mu(z)\lesssim\int\frac{\ell_{k}}{({\varepsilon}+|x-z|)^{n+1}}\,d\mu(z)\lesssim{\mathcal{P}}_{\mu}(Q).$
@@
$\int_{2B_{0}}|{\mathcal{R}}\sigma_{k}-m_{\sigma_{k},2B_{0}}({\mathcal{R}}\sigma_{k})|^{2}\,d\sigma_{k}$ $\leq\int_{2B_{0}}|{\mathcal{R}}\sigma_{k}-\rho|^{2}\,d\sigma_{k}$
$\leq 2\int_{2B_{0}}|{\widetilde{{\mathcal{R}}}}_{\ell_{k}}\mu-\rho|^{2}\,d\mu+C\int_{2B_{0}}{\mathcal{P}}_{\mu}(B(x,\ell_{k}))^{2}\,d\mu(x).$
@@
$\lim_{k\to\infty}\int_{2B_{0}}|{\widetilde{{\mathcal{R}}}}_{\ell_{k}}\mu-\rho|^{2}\,d\mu=\int_{2B_{0}}|{\mathcal{R}}\mu(x)-\rho|^{2}\,d\mu=\int_{2B_{0}}|{\mathcal{R}}\mu-m_{\mu,2B_{0}}({\mathcal{R}}\mu)|^{2}\,d\mu.$
@@ 5. The main lemma
Hence there is a compact set $E\subset B$ such that $\mu(E)\geq\mu(B)/4$ and $M_{n}(\chi_{B}\mu)(x)\leq 2C_{2}\,\Theta_{\mu}(B)$ for all $x\in E$. So the measure
@@
    (5.3) $|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))(x)\leq 16\lambda\,\Theta_{\mu}(B_{1})^{p}.$
@@
The fact that $F(a)\leq F(1)\leq 4\lambda\sigma_{p}(B_{1})$ also implies that
@@
To prove (c), for any ball $B$ centered in $R_{0}\cap\operatorname{supp}\nu_{a}$ and any $t\in(0,1)$, we consider the function
@@
Notice that $g$ is differentiable in $[0,1)$ and $g(t)\geq F(a_{t})\geq F(a)=g(0)$, by the minimizing property of $a$. Hence $g^{\prime}(0+)\geq 0$. Computing, we get
@@
Let $\Omega\subset{\mathbb{R}}^{n+1}$ be a bounded open set, let $\xi\in\partial\Omega$, and let $r_{0}>0$ and $r_{1}=8^{-N}r_{0}$, for some $N\geq 1$. Let $u$ be a nonnegative function which is continuous in $\overline{B(\xi,r_{0})\cap\Omega}$ and subharmonic in $B(\xi,r_{0})\cap\Omega$, and vanishes on $B(\xi,r_{0})\cap\partial\Omega$. Suppose that there are $c_{3}>0$ and a subset $J\subset\{1,\ldots,N\}$ with $J\geq\kappa N$, for some $\kappa>0$, such that
@@
Let $\mu$ and $B_{0}$ satisfy the assumptions of Main Lemma 5.1 and let $R_{0}$, $c_{R_{0}}$ be as in Lemma 5.3. Suppose that, for some $p\in(1,1+\frac{1}{n})$ and some $\lambda\in(0,1)$,
@@
$|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))(x)\leq\big[16\,\lambda+C(\alpha,\eta,\tau)\,\delta_{1}^{\kappa_{1}}\big]\,\Theta_{\mu}(B_{1})^{p}$
@@
$f(x)=\max\big(0,\,|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))(x)-16\lambda\,\Theta_{\mu}(B_{1})^{p}\big).$
@@
$\quad+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))(x).$
@@
(5.13) $|{\mathcal{R}}\nu(x)-c_{R_{0}}|^{p}+p\,{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))(x)\leq\gamma\,\Theta_{\mu}(B_{1})^{p}$
@@
$\leq C\gamma\Theta_{\mu}(B_{1})^{p}\int|\nabla{\varphi}|\,d{\mathcal{L}}^{n+1}-p\int{\mathcal{R}}_{\nu}^{*}(\chi_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))\,|\nabla{\varphi}|\,d{\mathcal{L}}^{n+1}$
$\leq C\gamma\Theta_{\mu}(B_{1})^{p}\,\mu(B_{1})-p\int_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))\,{\mathcal{R}}_{{\mathcal{L}}^{n+1}}(|\nabla{\varphi}|)\,d\nu.$
To bound the last integral on the right hand side we use the fact that, for any $x\in{\mathbb{R}}^{n+1}$,
@@
$\bigg|\int_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}$ $|^{p-2}({\mathcal{R}}\nu-c_{R_{0}}))\,{\mathcal{R}}_{{\mathcal{L}}^{n+1}}(|\nabla{\varphi}|)\,d\nu\bigg|$
$\lesssim\Theta_{\mu}(B_{1})\int_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p-1}\,d\nu$
$\leq\Theta_{\mu}(B_{1})\left(\int_{R_{0}}|{\mathcal{R}}\nu-c_{R_{0}}|^{p}\,d\nu\right)^{1/p^{\prime}}\nu(R_{0})^{1/p}$